# Mitigating Feature Gap for Adversarial Robustness by Feature Disentanglement

### v1

@@ 3.4 Adversarial Fine-tuning Based on Alignment
0:  Training set $\mathcal{D}$, The $\mathcal{L}_{1}$ weight $\alpha$, The $\mathcal{L}_{2}$ weight $\beta$, Number of epochs $T$, Learning rate $\eta$, Inner maximization $\mathcal{A}$, Naturally pre-trained model parameters $\theta^{P}$, Model parameters except linear block $D_{2}$ parameters $\theta_{1}$, Linear block $D_{2}$ parameters $\theta_{2}$.
@@
        Calculate loss $\mathcal{L}_{1}$ by Formula 2;
@@
        Calculate loss $\mathcal{L}_{2}$ by Formula 3;
@@
        Calculate loss $\mathcal{L}_{3}$ by Formula 4;
@@
4 Experiments
This section presents experiments with AFD. We first introduce our experiment setting. We further make comparisons on multiple architectures and datasets to show our superiority. Then we conduct an ablation study to show the functions of the disentangling module and alignment. Finally, we conduct an empirical analysis to support our hypothesis.
4.1 Setting
Datasets We conduct our experiments on three benchmark datasets including on CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), Tiny-ImageNet (Deng et al., 2009). CIFAR-10 dataset contains 60,000 color images having a size of $32\times 32$ in 10 classes, with 50,000 training and 10,000 test images. CIFAR-100 dataset contains 50,000 training and 10,000 test images in 100 classes. Tiny-ImageNet dataset contains 100000 images of 200 classes (500 for each class) downsized to 64×64 colored images. Each class has 500 training images, 50 validation images and 50 test images. We only use training and validation images of Tiny-ImageNet in experiments.
Baselines To make a comprehensive comparison, we have employed various techniques to demonstrate our superiority. Our primary baselines are AFT methods including vanilla AFT (vAFT) (Jeddi et al., 2020), AFT+RGKD (AFKD) (Suzuki et al., 2023), and ARREST (Suzuki et al., 2023). All of them are built upon naturally pre-trained models. Additionally, we compare our AFD with state-of-the-art AFT methods such as RiFT (Zhu et al., 2023) and TWINS (Liu et al., 2023), which rely on adversarially pre-trained models. Besides, we select several advanced fast adversarial training methods for comparison, including FreeAT (Shafahi et al., 2019), FGSM-GA (Andriushchenko & Flammarion, 2020), ATAS (Huang et al., 2023), since they are also low-cost techniques. Moreover, we compare with baseline methods of adversarial training including PGD-AT (Madry et al., 2017), TRADES (Zhang et al., 2019) to show our great advance in robustness.
Optimization Details The optimizer used in all the experiments is Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9. We naturally pre-train all the models with a batch size of 128, training epochs of 100, and a weight decay of ${5.0\times 10^{-4}}$. The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90\}$, following Zhang et al. (2019). For AT methods, we train all the models with a batch size of 128, training epochs of 120, and a weight decay of ${2.0\times 10^{-4}}$. The learning rate starts at 0.1 and then decays by $\times$0.1 with transition epochs $\{75,90,100\}$, following Wang et al. (2020). For our AFD, we fine-tune the pre-trained models with a batch size of 128, an initial learning rate of 0.0025, a weight decay of ${5.0\times 10^{-4}}$, and training epochs of 20. The learning rate is changed to 0.0020 at the 11th epoch and is divided by half every two epochs thereafter, as Suzuki et al. (2023).
@@ 4.1 Setting
4.2 Main Results against White-Box Attack
To evaluate AFD’s effectiveness, we conducted comparative experiments with other methods on CIFAR-10, CIFAR-100, and Tiny-ImageNet. We also make additional experiments on ResNet50 to fairly compare with TWINS.
CIFAR-10 and CIFAR-100 As shown in Tables 1, methods above the double line denote AT and FAT methods, and methods under the double line denote AFT methods. It shows that among AFT methods, AFD exhibits the best robust performance and ARREST has the second best robust performance on ResNet18. Robust accuracies of AFD against AutoAttack (AA) surpass those of ARREST by 1.27% and 2.58% on CIFAR-10 and CIFAR-100, respectively. As indicated in Table 2, AFD exhibits the best robust performance and AFKT has the second best robust performance among all the methods on WideResNet-34-10. AFD achieves superior robust accuracies against AA compared to AFKD, with improvements of 2.05% and 2.34% on CIFAR-10 and CIFAR-100, respectively. In particular, RiFT doesn’t show great robust performances, since it is fine-tuned with natural data to enhance out-of-distribution performance.
Besides, AFD with fewer training epochs outperforms AT methods. AFD surpasses PGD-AT in all respects. TRADES achieves a larger robust accuracy against AA on CIFAR-10 on ResNet18 than AFD by 1.02%. However, the situation is reversed on CIFAR-100 or on WideResNet-34-10, where AFD outperform TRADES against AA by 0.43%, 2.40%, and 2.37%. This can be attribute to the superiority of AFD and the robust over-fitting in AT. Moreover, FAT methods distinctly lag behind AFD in both natural and robust performances. Despite the advanced robust performance of AFD, its natural accuracy doesn’t surpass state-of-the-art AFT methods (ARREST). We make further discussion in the Section 4.3. In summary, AFD demonstrates superb robustness compared to AT, FAT and AFT methods.
Table 1: Quantitative evaluations of AFD and other methods for CIFAR-10 and CIFAR-100 on ResNet18.
Methods CIFAR-10 CIFAR-100
Clean PGD AA Clean PGD AA
PGD-AT 84.43 48.82 43.68 59.09 24.05 20.67
TRADES 82.37 53.27 48.52 55.47 28.12 23.49
${}^{{\dagger}}$ 78.37 40.90 36.00 50.56 19.57 15.09
${}^{{\dagger}}$ 80.10 49.14 43.44 50.61 24.48 19.42
${}^{{\dagger}}$ 81.22 50.03 45.38 55.49 27.68 22.62
vAFT 83.93 51.40 45.48 60.83 26.55 21.96
RiFT 84.49 49.24 43.66 59.51 24.01 20.48
AFKD 84.71 51.32 46.09 65.54 28.09 22.92
ARREST 85.71 51.23 46.23 67.00 26.89 21.34
AFD 84.66 52.63 47.50 62.94 29.27 23.92
Table 2: The performance of AFD and other methods on CIFAR-10 and CIFAR-100 on WideResNet-34-10.
Methods CIFAR-10 CIFAR-100
Clean PGD AA Clean PGD AA
PGD-AT 86.35 49.57 45.81 60.30 25.14 22.65
TRADES 85.68 52.53 49.22 57.67 28.26 25.25
vAFT 87.42 53.57 48.49 64.92 27.19 23.59
RiFT 84.49 49.24 45.83 60.59 25.57 22.67
AFKD 88.54 54.86 49.57 68.50 30.12 25.28
ARREST 88.42 54.16 49.45 69.52 29.47 24.55
AFD 88.08 56.12 51.62 68.51 32.63 27.62
@@ 4.2 Main Results against White-Box Attack
Tiny-ImageNet We further conduct comparative experiments on a larger dataset, Tiny-ImageNet. As shown in Table 3, on ResNet18, AFD consistently outperforms all the methods in terms of robust performance and the robust performance of TRADES ranks second. The robust accuracy of AFD against PGD exceeds that of TRADES by 1.59% and its robust accuracy against AA outpaces that of TRADES by 1.58%. Moreover, on WideResNet-34-10, AFD keeps the best robust performance and AFKD exhibits the second best robust performance. The robust accuracy of AFD against PGD on WideResNet-34-10 surpasses that of AFKD by 2.56% and its robust accuracy against AA on WideResNet-34-10 is higher than that of AFKD by 2.76%. These findings indicate the superiority of AFD in terms of adversarial robustness on large datasets.
Table 3: The performances of AFD and other methods on Tiny-ImageNet with ResNet18 and WideResNet-34-10 models.
Methods ResNet18 WideResNet-34-10
Clean PGD AA Clean PGD AA
PGD-AT 52.12 19.38 16.79 54.24 19.03 16.51
TRADES 49.28 22.59 17.06 49.22 23.33 18.51
vAFT 52.75 21.29 16.76 57.17 22.24 18.02
RiFT 52.24 20.89 17.05 55.09 19.51 16.79
AFKD 57.17 22.29 16.68 60.92 24.27 19.01
ARREST 60.36 19.28 12.58 65.07 22.73 16.67
AFD 55.58 24.18 18.64 59.58 26.83 21.77
ResNet50 We follow Liu et al. (2023) to make comparisons on ResNet50. Notice that TWINS employs a model adversarially pre-trained on ImageNet (Deng et al., 2009), and the number of its training periods is 60. We set the training epochs of AFD as 40. Results are shown in Table 4. Although TWINS benefits from prior knowledge of ImageNet and adversarial robustness of the pre-trained model, AFD still surpasses its robust performance on both CIFAR-10 and CIFAR-100. AFD achieves improvements of 1.54% and 1.56% under AA compared to TWINS, and its robust accuracies against PGD surpass those of TWINS by 1.88% and 2.52%. This demonstrates its superiority in terms of adversarial robustness on ResNet50.
Table 4: The performances of AFD and the state-of-the-art method TWINS on CIFAR-10 and CIFAR-100 on ResNet50.
Methods CIFAR-10 CIFAR-100
Clean PGD AA Clean PGD AA
${}^{{\dagger}}$ 89.77 52.24 48.46 69.48 28.52 23.47
${}^{{\dagger}}$ 91.95 52.46 49.02 72.12 29.12 25.72
AFD 86.31 54.34 50.56 65.48 31.64 27.28
In summary, these experiments indicate that AFD always outperforms all the baselines for different adversarial attacks, datasets, and network architectures in the aspect of robustness.
4.3 Ablation Study
Effect of Disentangler As mentioned in Section 3.3, the disentangler should disentangles adversarial features into the confused and unconfused features, and further eliminate the confused features. As shown in Table 5, vAFT with the disentangler (vAFT+D) exhibits superior robustness compared to vAFT by 1.28% against PGD and by 1.40% against AA, respectively. It shows the feature disentangler effectively eliminates the confused features that lead to wrong predictions. Moreover, the natural accuracy of AFT-D outperforms that of AFT by 0.70%. This could be attributed to the disentangling module removing the class-irrelevant features of natural samples such as background and shadow.
Effect of Alignment As shown in Table 5, vAFT with the alignment of features (vAFT+A) exhibits superior robustness compared to AFT against both PGD and AA by 0.39% and 0.72%, with a marginal increase in natural accuracy by 0.67%. It indicates that the alignment successfully assists to further remove the confused features.
Effect of Combination The combination of disentangler and alignment (AFD) has the best performance on both natural and robust accuracy, which shows two strategies are compatible and the combination can boost the positive effects of individual strategies. However, in Table 1 and other tables, AFD exhibits a lower natural accuracy than state-of-the-art AFT methods. This is because the unconfused features $f_{1}$ have not been trained with natural samples, which may lead to an accuracy-robust trade-off.
Methods vAFT vAFT+D vAFT+A AFD
Clean 83.93 84.63 84.61 84.66
PGD 51.40 52.68 51.71 52.63
AA 45.48 46.88 46.81 47.50
Figure 5: $L_{p}$ distances of features between natural and adversarial samples. The horizontal axis represents the training epoch.
Effect of Hyperparameter There are three hyperparameters in our AFD: $\alpha,\beta$ and $\gamma$. They control the weights of $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and the distance function in $\mathcal{L}_{3}$, respectively. We have conducted experiments with various combinations of weights on ResNet18 on CIFAR-10, as illustrated in Figure 4. There is a remarkable trade-off between accuracy and robustness when $\beta$ and $\gamma$ are growing. On the one hand, a larger value of $\beta$ indicates a larger exclusive force between the unconfused and confused features, leading to less non-robust features utilized in classification and better robustness. On the other hand, as $\gamma$ increases, the natural accuracy increases but robust accuracy drops. This occurs because more feature information of natural samples is transferred to the fine-tuned model through the feature alignment. However, the variations of both natural and robust accuracy are less than 0.60%, demonstrating that our AFD is insensitive to variations of hyperparameters.
4.4 Empirical Analysis of Disentangled Features
@@
We evaluate the performances of two types of disentangled features $f_{1}$, $f_{2}$ on various datasets and models. As shown in Table 6, the natural accuracies of $f_{2}$ are lower than those of $f_{1}$ by $4.53\%\sim 45.05\%$. It indicates the disentangler has successfully extracted and removed the confused features as $f_{2}$. High accuracies of $f_{1}$ indicates models benefit from eliminating the confused features. Interestingly, the disparity of accuracy on CIFAR-100 is smaller than that on CIFAR-10. We suppose that a dataset with more classes has more complex feature space, and it is more difficult to extract confused features by the linear disentangler. Maybe nonlinear disentanglers work better.
Table 6: The natural accuracies of the unconfused features $f_{1}$ and confused features $f_{2}$. RN and WRN are abbreviations of ResNet and WideResNet, respectively. $\Delta$ denotes the accuracy disparity.
@@
RN18 CIFAR-10 84.66 39.61 45.05
RN18 CIFAR-100 62.94 58.41 4.53
RN50 CIFAR-10 83.18 47.35 35.83
RN50 CIFAR-100 61.68 54.97 6.71
@@
Moreover, we analyze the feature distance between the unconfused features $f_{1}$ and natural features $f_{nat}$, as illustrated in Figure 5. In contrast to the growing trend of feature distance in existing AFT methods, the optimization trajectory of the feature distance in AFD exhibits a fast convergence trend. It implies that AFD has mitigated the gap between natural and adversarial features. Additionally, the magnitudes of the feature distance in AFD are smaller than other AFT methods, e.g., the $L_{\infty}$ distance of AFD (0.214) is notably less than those of ARREST and vanilla AFT (0.453, 0.236). These results demonstrates our disentanglement-based approach can enhance the adversarial robustness by alleviating the feature gap.
5 Conclusion

### v2

@@ Introduction
0:  Training set $\mathcal{D}$, Hyperparameters $\alpha,\beta$, Number of epochs $T$, Learning rate $\eta$, Naturally pre-trained model parameters $\theta^{P}$, Model parameters except linear block $D_{2}$ parameters $\theta_{1}$, Linear block $D_{2}$ parameters $\theta_{2}$.
@@
        Calculate loss $\mathcal{L}_{1}$ Formula 3;
@@
        Calculate loss $\mathcal{L}_{2}$ Formula 4;
@@
        Calculate loss $\mathcal{L}_{3}$ by Formula 5;
@@
Experiments
This section presents experiments with AFD. We first introduce our experiment setting. We further make comparisons on multiple architectures and datasets to show the effectiveness. Then we conduct an ablation study to show the functions of the disentangling module and alignment. Finally, we conduct an empirical analysis to support our hypothesis.
Setting
Datasets We conduct experiments on three benchmark datasets including CIFAR-10 and CIFAR-100 (Krizhevsky 2009), Tiny-ImageNet (Deng et al. 2009). CIFAR-10 dataset contains 60,000 color images having a size of $32\times 32$ in 10 classes, with 50,000 training and 10,000 test images. CIFAR-100 dataset contains 50,000 training and 10,000 test images in 100 classes. Tiny-ImageNet dataset contains 100000 images of 200 classes (500 for each class) downsized to 64×64 colored images. Each class has 500 training images, 50 validation images and 50 test images. We only use training and validation images of Tiny-ImageNet.
Baselines To make a comprehensive comparison, we have employed various techniques to demonstrate the effectiveness of our method AFD. Our primary baselines are AFT methods including vanilla AFT (vAFT) (Jeddi, Shafiee, and Wong 2020), AFT+RGKD (AFKD) (Suzuki et al. 2023), and ARREST (Suzuki et al. 2023). All of them are built upon naturally pre-trained models. Additionally, we compare our AFD with a state-of-the-art AFT method RiFT (Zhu et al. 2023), which relies on adversarially pre-trained and has powerful out-of-distribution performances. Besides, we select several advanced fast adversarial training methods, including FreeAT (Shafahi et al. 2019), FGSM-GA (Andriushchenko and Flammarion 2020), ATAS (Huang et al. 2023), since they are also low-cost techniques. Moreover, we compare with baseline methods of adversarial training including PGD-AT (Madry et al. 2017), TRADES (Zhang et al. 2019) to show our great advance in robustness.
Optimization Details The optimizer is Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9. We fine-tune the pre-trained models with a batch size of 128, a learning rate of 0.0025, a weight decay of ${5.0\times 10^{-4}}$, and training epochs of 20. The schedule of learning rate is the same as Suzuki et al. (2023). We utilize Exponential Moving Average (EMA) (Hunter 1986) to gain better parameters.
@@
Methods CIFAR-10 CIFAR-100 Tiny-ImageNet
Clean PGD AA Avg Clean PGD AA Avg Clean PGD AA Avg
PGD-AT 84.43 48.82 43.68 58.98 59.09 24.05 20.67 34.60 52.12 19.38 16.79 29.43
TRADES 82.37 53.27 48.52 61.39 55.47 28.12 23.49 35.69 49.28 22.59 17.06 29.64
Free-AT 78.37 40.90 36.00 51.76 50.56 19.57 15.09 28.41 - - - -
FGSM-GA 80.10 49.14 43.44 57.56 50.61 24.48 19.42 31.50 - - - -
ATAS 81.22 50.03 45.38 58.88 55.49 27.68 22.62 35.26 - - - -
vAFT 83.93 51.40 45.48 60.27 60.83 26.55 21.96 36.45 52.75 21.29 16.76 30.27
RiFT 84.49 49.24 43.66 59.13 59.51 24.01 20.48 34.67 52.24 20.89 17.05 30.06
AFKD 84.71 51.32 46.09 60.71 65.54 28.09 22.92 38.85 57.17 22.29 16.68 32.05
ARREST 85.71 51.23 46.23 61.05 67.00 26.89 21.34 38.41 60.36 19.28 12.58 30.74
AFD 84.88 52.70 47.40 61.66 65.69 28.56 23.05 39.10 57.41 23.63 17.73 32.92
Table 1: Quantitative evaluations of all methods on ResNet18.
Methods CIFAR-10 CIFAR-100 Tiny-ImageNet
Clean PGD AA Avg Clean PGD AA Avg Clean PGD AA Avg
PGD-AT 86.35 49.57 45.81 60.58 60.30 25.14 22.65 36.03 54.24 19.03 16.51 29.93
TRADES 85.68 52.53 49.22 62.48 57.67 28.26 25.25 37.06 49.22 23.33 18.51 30.35
vAFT 87.42 53.57 48.49 63.16 64.92 27.19 23.59 38.57 57.17 22.24 18.02 32.48
RiFT 84.49 49.24 45.83 59.85 60.59 25.57 22.67 36.28 55.09 19.51 16.79 30.46
AFKD 88.54 54.86 49.57 64.31 68.50 30.12 25.28 41.30 60.92 24.27 19.01 34.73
ARREST 88.42 54.16 49.45 64.01 69.52 29.47 24.55 41.18 65.07 22.73 16.67 34.82
AFD 88.08 56.12 51.62 65.27 68.51 32.63 27.62 42.92 63.89 25.73 19.13 36.25
Table 2: Quantitative evaluations of all methods on WideResNet-34-10.
Main Results
ResNet18 First, AFD effectively mitigates the trade-off between generalization and robustness. As shown in Table 1, AFD takes the lead in robust performance and exhibits the second-best clean accuracy among AFT methods. It gains less clean accuracies compared to ARREST by 1.70%, but its robust accuracies against PGD-20 and AA outperform ARREST by 2.50% and 2.68%. The improvement in robustness is evidently larger than the drop in clean accuracy. This result indicates the positive effect of feature disentanglement on the accuracy-robustness trade-off.
Moreover, AFD exceeds almost all the AT and FAT methods. Although the powerful AT baseline TRADES outperforms AFD in the robust accuracies on CIFAR-10 and CIFAR-100 by 0.56%, it surpasses TRADES in clean and average accuracy by 6.365% and 1.84%. Besides, it surpasses TRADES on Tiny-ImageNet in all metrics. The result implies that efficient AFD can achieve comparable performances with AT.
WideResNet-34-10 Second, AFD takes further advantages on large architectures. As shown in Table 2, AFD maintains the best robust and average accuracies among all the approaches. It exceeds the second-best baselines by approximately 1.62% and 1.34%. The enhancement of robustness further enlarges compared to that on ResNet18. In addition, the natural generalization of AFD always ranks second and is less than the best clean accuracy by 0.88% on average. The discrepancy of clean accuracies between AFD and the best one is smaller than that on ResNet18 (1.69%). Furthermore, AFD outperforms AT baselines in all metrics. These demonstrate that AFD achieves a better trade-off between generalization and robustness on larger architectures.
It’s worth noting that AFD doesn’t exhibit optimal natural generalization among AFT methods. It is attritued to merely adversarial features purified by disentangling modules. Intuitively, natural features can also be boosted by disentanglement mechanism. Nevertheless, it will result in additional computing cost. Besides, current AFD has remained the second-best clean accuracy among various methods.
Summary These experiments indicate that AFD always outperforms all the baselines for different attacks, datasets, and architectures in the aspect of overall performance. It proves the effectiveness of our AFD.
Ablation Study
Disentangler and Alignment As shown in Table 3, vanilla AFT with the disentangler (vAFT+D) exhibits superior performances compared to vanilla AFT by 0.70%, 1.28%, and 1.40%, respectively. It shows that the disentangler effectively eliminates the confused features that lead to wrong predictions. Besides, vanilla AFT with the alignment of features (vAFT+A) outperforms vanilla AFT by 0.68%, 0.31%, and 1.33%. It indicates that the alignment successfully contributes to removing the confused features. Moreover, the combination of disentangler and alignment (AFD) has the best performance on both natural and robust accuracy, which shows two strategies are compatible and the combination can boost the positive effects of individual strategies.
Methods vAFT vAFT+D vAFT+A AFD
Clean 83.93 84.63 84.61 84.88
PGD 51.40 52.68 51.71 52.70
AA 45.48 46.88 46.81 47.40
Hyperparameter There are three hyperparameters in our AFD: $\alpha,\beta$ and $\gamma$. As illustrated in Figure 4, there is a remarkable trade-off between accuracy and robustness when $\beta$ and $\gamma$ are growing. On the one hand, a larger value of $\beta$ indicates a larger exclusive force between the intrinsic and confused features, leading to less non-robust features utilized in classification and better robustness. On the other hand, as $\gamma$ increases, the natural accuracy increases but robust accuracy drops. This occurs because more feature information of natural samples is transferred to the fine-tuned model through the feature alignment. Besides, the variations of both natural and robust accuracies are less than 0.60%, demonstrating that our AFD is insensitive to variations of hyperparameters. This conclusion is also true among different datasets.
Figure 3: t-SNE visualization on CIFAR-10 on ResNet18.
@@
Empirical Analysis of Disentangled Features
we analyze the feature distance between the intrinsic features $f_{1}$ and natural features $f_{nat}$, as shown in Figure 1 (a). In contrast to the growing trend of feature distance in existing AFT methods, the optimization trajectory of the feature distance in AFD exhibits a fast convergence trend. It implies that AFD has mitigated the gap between natural and adversarial features. Additionally, the magnitudes of the feature distance in AFD are smaller than other AFT methods, e.g., the $L_{\infty}$ distance of AFD (0.212) is notably less than those of ARREST and vanilla AFT (0.453, 0.236). These results demonstrate our AFD can enhance the adversarial robustness by alleviating the feature gap.
@@
Moreover, we evaluate the performances of two types of disentangled features $f_{1}$, $f_{2}$ on various datasets and models. As shown in Table 4, the natural accuracies of $f_{2}$ are lower than those of $f_{1}$ by $5.67\%\sim 51.04\%$. It indicates the disentangler has successfully extracted and removed the confused features as $f_{2}$. High accuracies of $f_{1}$ indicate models benefit from eliminating the confused features. Interestingly, the disparity of accuracies on CIFAR-100 is smaller than those on CIFAR-10 and Tiny-ImageNet. We suppose that it is challenging to learn clear feature space from a small-resolution dataset with numerous classes, which made the linear disentangler difficult to extract confused features.
We make further feature visualization by t-SNE (Van der Maaten and Hinton 2008). We take the fine-tuned model by AFD. The we acquire features of adversarial samples crafted by PGD-20. We visualizes two types of features: the features before disentanglement and the intrinsic features $f_{1}$. As illustrated in Figure 3, features after disentanglement has a distincter class boundary than features before disentanglement, indicating the effectiveness of disentanglement.
@@
RN18 CIFAR-10 84.88 39.05 45.83
RN18 CIFAR-100 65.69 58.17 7.52
RN18 Tiny-ImageNet 57.41 48.78 8.63
RN50 CIFAR-10 87.01 35.97 51.04
RN50 CIFAR-100 67.45 56.17 11.28
@@
WRN-34-10 Tiny-ImageNet 63.89 55.15 8.74
Table 4: The natural accuracies of the intrinsic features $f_{1}$ and confused features $f_{2}$. RN and WRN are abbreviations of ResNet and WideResNet, respectively. $\Delta$ denotes the accuracy disparity.
Training Time and Model Parameter
[…]
The vanilla ResNet18 and WideResNet-34-10 have parameters of 11.2M and 46.2M. In the training process, the additional module has a weight matrix that doesn’t depend on the input size. It is $2*512*512=0.5$M (4.46% increase) on ResNet18 and $2*1024*1024=2$M (4.33% increase) on WideResNet-34-10, which is acceptable for better training.
Method Epoch ${\rm Time_{one}}$ ${\rm Time_{total}}$
Standard 100 6.0 590
PGD-AT 120 37 4440
PGD-AT - - 4428
TARDES 120 53 6360
Free-AT 10 119 1188
FGSM-GA 30 68 2052
ATAS 30 36 1080
vanilla AFT 20 37 1330
RiFT 10 25 4690
AFKD 20 40 1390
ARREST 20 36 1310
AFD 20 46 1510
Table 5: Training time (second) of various methods on ResNet18 on CIFAR10. We show the time for one epoch (${\rm Time_{one}}$) and the expected values of the total time (${\rm Time_{total}}$). ‘Standard’ denotes the natural pre-training.
Limitation and futural work
Subject to computation resources, we haven’t conducted experiments on ImageNet-1k. We will explore effective adversarial fine-tuning approaches on large-resolution datasets in the future. Besides, we observe that the increasing gap of features also appears in adversarial training. It remains a lot for further studies.
Conclusion